# Unravelling the asphericities in the explosion and multi-faceted circumstellar matter of SN 2023ixf

### v1

@@ 5.2 Photospheric Phase: 8 - 80 d
Beginning 16 d, we observe a more intricate and complex multi-peaked profile of with a persistent high-velocity absorption feature, which we extensively discuss in Section 5.3. In addition to the distinguishable Balmer lines in the early spectra, several distinct features start to appear 16 d onwards, particularly Ca II H&K and He I 5876 Å. As the photosphere recedes further into the ejecta, we detect the appearance of various lines of Fe II, Sc II and Ba II, and the Ca II Triplet beginning 25 d. These metal features are visibly developed around 36 d. The broad Na ID starts to emerge at the location of He I 5876, evident from the slightly broadened absorption trough at around 40 d (continuing up to 54 d). This is further indicative of the cooling of the ejecta and is suggestive of an ejecta that is cooling a lot slower than typical SNe II as the appearance of Na ID, is usually observed around 30 d (Gutiérrez et al., 2017a). Towards the end of the plateau drop, i.e. around 70 d, we also begin to observe O I 7774.
@@
In the top-right and bottom-right panels of Figure 5, we compare the spectrum of SN 2023ixf with other SNe II spectra encompassing a sample of SNe IIL and short plateau SNe i.e. SN 2013by (Valenti et al., 2015); SN 2014G (Terreran et al., 2016); SN 2018gj (Teja et al., 2023b); SN 2020jfo (Teja et al., 2022); SN 2020pni (Terreran et al., 2022); SN 2020tlf (Jacobson-Galán et al., 2022). Photospheric features develop pretty late in SN 2023ixf compared to other Type IIL and short-plateau SNe II, significantly later than normal SNe II. In the early plateau phase ( 25 d), the spectra blueward of H is generally dominated by metallic features as in the case for SN 2014G, SN 2020jfo and SN 2020pni; however, the features are relatively underdeveloped in SN 2023ixf and SN 2018gj. The line strengths appear much weaker in SN 2023ixf than in spectral lines of other SNe. This also plausibly hints at the ejecta being still hot and/or metal-poor compared to normal SNe II. In addition, the P-Cygni absorption during the early photospheric phase in SN 2023ixf is a lot weaker than short-plateau SNe 2018gj and 2020jfo; however, it is similar to SNe 2014G and 2020pni. The spectral features are likely less prominent due to the luminosity from interaction enhancing the ejecta’s temperature (and ionization), leading to an ionization wave penetrating the ejecta inwards from the cold, dense shell (CDS, Chevalier & Fransson, 1994). CDS arises from the higher density of the inner CSM (, see Section 7) resulting from the cooling of shocked areas during the early phases.
This is also in concordance with Type IIL SNe, which tend to show weaker P-Cygni absorption troughs since these SNe likely have more significant interaction with CSM than typical Type IIP SNe (Gutiérrez et al., 2014; Polshaw et al., 2016). Upon examination of the spectral comparison during the late-plateau phase, we observe higher line blending in SN 2023ixf due to its higher photospheric velocity (see Section 5.4) during the late-plateau phase in comparison to all other SNe II in our comparison except that of SN 2014G.
@@ 5.3 Evolution of H : High-Velocity Absorption and Clumpy CSM
The emission peak of has a blueshifted offset by as much as 3,000 at 32 d but evolves towards zero rest velocity by the end of the plateau. This is a result of the steep density profile of the ejecta leading to a higher occultation of the receding portion of the ejecta (Anderson et al., 2014b). The blueshifted offset of the emission peak of H (at 30 d) typically correlates with the decline rate during the plateau phase and the peak luminosity of SNe II (Anderson et al., 2014b). The steep decline rate of SN 2023ixf ( is in agreement with its large offset of 3,000 indicating that the ejecta mass is smaller in SN 2023ixf leading to its shorter plateau. Although this effect is commonly seen in SNe II, this effect is more pronounced in only a handful of events, e.g., SN 2014G (Terreran et al., 2016) and SN 2018gj (Teja et al., 2023b), where the emission peak stays blue-shifted late into the nebular phase.
@@ 5.4 Line velocity Evolution
We compare the velocity evolution of Fe II 5169 for SN 2023ixf with other SNe II in Figure 7. We estimated the line velocity evolution from the blue-shifted absorption trough of their line profiles in the redshift-corrected spectral sequence of SN 2023ixf. We adopt the Fe II 5169 velocities as the photospheric velocity since it forms closest to the photosphere (Dessart & Hillier, 2005) and is the least blended among the iron lines seen in our spectral sequence. The photospheric radius estimated from Fe II absorption trough closely mirrors the blackbody radius as shown in Figure 4. During the mid-plateau phase ( 53 d), the Fe II 5169 velocity inferred for SN 2023ixf is 4350 . This positions it at the higher end of the 1- range when compared with the mean velocity of Fe II 5169 for an extensive collection of SNe II by Gutiérrez et al. (2017a, 3537 851 km/s). The trend continues onto the early nebular phase ( 116 d), where SN 2023ixf displays a Fe II 5169 velocity of 3330 compared to the mean value of 2451 679 from Gutiérrez et al. (2017a). The higher photospheric velocity is in agreement with the luminosity-velocity correlation of homologously expanding recombination front of hydrogen (Kasen & Woosley, 2009) since SN 2023ixf is brighter during the mid-plateau phase compared to a normal Type II SN (Anderson et al., 2014a). Additionally, the CSM interaction in SN 2023ixf could also drive the elements at outer/faster regions of the ejecta to be reionized and recombined, leading to a higher estimate of photospheric velocity (Andrews et al., 2019).
@@ 5.5 Early Nebular Phase 90 d
The nebular-phase spectral sequence of SN 2023ixf is shown in the left panel of Figure 8. The nebular spectra of SN 2023ixf displays prominent emission features of H , Mg I] 4571, H , O I 5577, Na I D / He I, [O I] 6300, 6364, H , [Ca II] 7291, 7324 and the Ca II NIR triplet with a flat continuum, typical of SNe II. We observe an asymmetry in the line profile of certain emission features during the nebular phase. We observe a dual-peaked axisymmetric888The separation between the two components is less than that of the two components of the O I doublet profile of the [O I] doublet in SN 2023ixf during the early nebular phase in Figure 9. We also observe an apparent redshifted excess in H and [Ca II] 7291, 7324 at around +5,000 possibly indicating asymmetries in the ejecta. However, the H and [Ca II] showed a single peak symmetric profile with visible signs of redward attenuation as the SN progressed into the nebular phase.
@@
The emergence of an asymmetric emission profile of H and [Ca II] arising due to the attenuation of the red-ward emission from the receding portions of the ejecta is first noticed in our spectral sequence beginning 125 d in Figure 9. This indicates the onset of dust formation and was first noticed in SN 1987A (Lucy et al., 1989). We also observe that the red-blue asymmetry increases as the SN evolves into the nebular phase, indicating an increased dust formation with time (Bevan et al., 2019). The early signatures of dust suggest its formation is happening inside the CDS since the SN ejecta during the 125 – 140 d is too warm for the condensation for molecules (Kozasa et al., 1991). The flash-ionization features in the early spectral sequence, and the steep rise in early UV light curves of SN 2023ixf conclusively indicated the presence of a dense CSM. As the shockwave from the SN encountered the denser CSM, it decelerated, compressing the material and increasing the density within the shocked CSM. Radiative cooling then facilitated the emission of photons, aiding in its cooling and forming a distinctive CDS (Chugai, 2009) in the denser CSM. The CDS enables an additional pathway for dust formation in interacting SNe II (Rho et al., 2018). In addition, the clumpiness within the extended CSM encompassing SN 2023ixf facilitates the formation of additional CDS (in addition to its formation in the dense CSM), consequently enhancing molecule formation and eventually forming dust (Inserra et al., 2011). Since nebular phase H arises from the inner ejecta, it wouldn’t show wavelength-dependent attenuation if the dust is formed in the outer CDS. This suggests that the regions where Balmer lines form and dust formation occurs essentially overlap, indicating thorough mixing of the CDS into the inner ejecta following a significant episode of CSM interaction (Bevan et al., 2019).
@@
We also see evidence of flattening in the -band light curve of SN 2023ixf beyond 125 d in Figure 2 evolving at 1.3 0.1 against the relatively consistent decline of 1.8 0.1 in the and bands. This indicates that the continuum luminosity in NIR is evolving steadily; however, the -band light curve is evolving rather slowly due to the emission from CO overtone around 2.3m. Although we don’t have spectral conformation in NIR, Type II SN 2017eaw (Rho et al., 2018) showed the presence of the first overtone of CO as early as 124 d and also showed flattening in their -band light curves. This strengthens our inference for indirect conformation of molecular CO formation and eventually dust in the case of SN 2023ixf.
@@ 6 Polarimetric Analysis
Figure 10 presents spectropolarimetry observations of SN 2023ixf during the early plateau phase. Due to the low signal-to-noise ratio of the individual epochs, we averaged out the first three epochs (June 14, 15, 16) and the following two epochs (June 18, 19) into two spectropolarimetric observations represented by their average phase of observations. The polarization is expressed as a sum of the and stokes parameters, P , and the position angle on the sky is given by the PA = 0.5 .
@@ 6.1 Estimating Interstellar Polarization
To analyze the intrinsic polarization of SN 2023ixf, it’s crucial to quantify the Interstellar Polarization (ISP) originating from the irregularly shaped dust particles present across the line of sight in both the Milky Way and the host galaxy M 101. To estimate an upper limit on the ISP, we assume a Serkowski-Galactic type ISP (Serkowski et al., 1975) for M 101 by assuming a similar size and dust composition as the Milky Way. Using the extinction values along the line-of-sight of SN 2023ixf in Section 2, the upper limit on the ISP from Serkowski et al. (1975) is 9 = 0.35%. Since the three peaks in our imaging polarization evolution of SN 2023ixf are all higher than the maximum limit on ISP, we reasonably conclude that the observed polarization characteristics originate from the SN itself.
@@ 6.2 Polarization in the Stokes Q-U Plane and Comparison with other SNe II
In Figure 11, we present ISP-corrected imaging polarization evolution of SN 2023ixf in -band spanning 1.4 d until 140 d. Figure 11 shows 3 distinct peaks in polarization labeled as Peak A [, ] at 1.4 d, Peak B [ at 6.4 d, ] and Peak C [, ] at 79.2 d. The observed stokes parameters ( and ) in -band are shown in Figure 12. Our first imaging polarimetric observation of SN 2023ixf was conducted at 1.4 d, corresponding to Peak A, and is associated with the flash-ionized phase due to the interaction of the SN ejecta with an asymmetric, compact, and dense CSM. This is similar to the value of continuum polarization inferred using spectropolarimetric data by Vasylyev et al. (2023). However, we observe a decrease in the polarization signature observed from the dense CSM on the 2nd epoch, i.e., 2.5 d, which is in stark contrast to the rising luminosity (see a six-fold rise in -band luminosity). However, the continuum polarization reported by Vasylyev et al. (2023) at the same epoch, after correction for ISP, indicates a slight polarisation increase from the first epoch, consistent with rising luminosity driven by interaction. The discrepancy could likely be due to the increased line depolarization of H, leading to a decrease in -band polarization as the PA remains the same.
@@
At the next epoch on 4.4 d, we see a change in position angle to 55 which later evolves to the polarization peak B at 6.4 d. The polarization peak B lies towards the end of the flash phase transitioning to the early photospheric phase, which shows interaction signatures in the form of intermediate-width P-Cygni profiles (Smith et al., 2023) and close to the emergence of the broad Balmer features at 7 d (see Section 5.2). Polarization at Peak B likely arises from interaction with extended CSM and/or an aspherical shock structure of SN 2023ixf. Following Peak B, we see a decline in polarization across the early plateau phase ( 22 d); however, the PA continues evolving slowly. There is slight disagreement with the continuum position angle reported by Vasylyev et al. (2023) likely due to the depolarization of the emission peak of the P-Cygni profile of H. We do observe a revival in the polarization evolution at 28 d with the position angle evolving from 100 to 150 . We have a 40-day gap before our next observations at 70 d, before the end of the plateau phase. The observed polarization evolution of SN 2023ixf after 70 d demonstrated a rise towards Peak C in polarization at 79.2 d towards the end of the plateau phase. Since the innermost layers of the ejecta ( 3000 ) are revealed as a result of the entire hydrogen envelope becoming optically thin, this suggests that the electron-scattering atmosphere of the inner core shows a substantial departure from spherical symmetry. The polarization dropped slowly across the early nebular phase, whereas the PA increased gradually. SN 2023ixf shows two preferred dominant axes: a) along the CSM-outer ejecta, where the SN evolved during the early phase until the epoch of polarization minimum on 21.3 d, and b) along the inner ejecta/core, where the SN evolved during the late-plateau and early nebular phase, which is inferred from its polarization evolution.
@@
In a recent study, Nagao et al. (2024) compared the intrinsic polarization properties of 15 SNe II. Broadly, it was inferred that the degree of polarization rises significantly () in the transition/nebular phase; however, it is usually lower during the photospheric phase (). Anomalous cases exist in SN 1987A (Jeffery, 1989) and SN 2021yja (Vasylyev et al., 2024), where a significant polarization is observed during the early plateau phase, gradually declining towards the end of the plateau phase. This is likely driven by electron scattering from interaction with an aspherical CSM, which is analogous to the early continuum polarization (Peak A) in SN 2023ixf. However, we also see a secondary peak (Peak B) in polarization in SN 2023ixf, which could possibly arise from interaction with a clumpy low-density extended CSM and/or an aspherical shock structure (see Section 8.4 for further discussions). Towards the end of the plateau phase, the swift increase in polarization during the late-plateau phase and steady decrease during the early nebular phase is similar to that seen in several SNe II like SN 1987A (Jeffery, 1989), SN 2004dj (Leonard et al., 2006) and SN 2013ej (Kumar et al., 2016). Such a polarization evolution is often interpreted as arising from the asymmetries in the inner He-core, and the trend seen is similar in SN 2023ixf. Type IIL SN 2017ahn, which showed early flash-ionization features similar to SN 2023ixf, showed low continuum polarization during the late-plateau phase photospheric and nebular phases, indicating that it might have been asymmetric; however, its viewing angle was lying across its polar axis (Nagao et al., 2021). However, it is clear from the observations of SN 2023ixf that the inclination differs from the axis of symmetry.

### v2

@@ 5.2 Photospheric Phase: 8 - 80 d
Beginning 16 d, we observe a more intricate and complex multi-peaked profile of with a persistent high-velocity absorption feature, which we extensively discuss in Section 5.3. In addition to the distinguishable Balmer lines in the early spectra, several distinct features start to appear 16 d onwards, particularly Ca II H&K and He I 5876 Å. As the photosphere recedes further into the ejecta, we detect the appearance of various lines of Fe II, Sc II and Ba II, and the Ca II Triplet beginning 25 d. These metal features are visibly developed around 36 d. The broad Na ID starts to emerge at the location of He I 5876, evident from the slightly broadened absorption trough at around 40 d (continuing up to 54 d). This is further indicative of the cooling of the ejecta (Gutiérrez et al., 2017a) and is suggestive of an ejecta that is cooling a lot slower than typical SNe II as the appearance of Na ID, is usually observed around 30 d (Gutiérrez et al., 2017a). Towards the end of the plateau drop, i.e. around 70 d, we also begin to observe O I 7774.
@@
In the top-right and bottom-right panels of Figure 5, we compare the spectrum of SN 2023ixf with other SNe II spectra encompassing a sample of SNe IIL and short plateau SNe i.e. SN 2013by (Valenti et al., 2015); SN 2014G (Terreran et al., 2016); SN 2018gj (Teja et al., 2023b); SN 2020jfo (Teja et al., 2022); SN 2020pni (Terreran et al., 2022); SN 2020tlf (Jacobson-Galán et al., 2022). Photospheric features develop pretty late in SN 2023ixf compared to other Type IIL and short-plateau SNe II, significantly later than normal SNe II. In the early plateau phase ( 25 d), the spectra blueward of H is generally dominated by metallic features as in the case for SN 2014G, SN 2020jfo and SN 2020pni; however, the features are relatively underdeveloped in SN 2023ixf and SN 2018gj. The line strengths of absorption features appear much weaker in SN 2023ixf than in spectral lines of other SNe. This plausibly hints at the ejecta being still hot and/or metal-poor compared to normal SNe II. We also observe that SN 2023ixf exhibits a weaker P-Cygni absorption during its early photospheric phase in tandem with short-plateau SNe from (Hiramatsu et al., 2021b) but not with short-plateau SNe 2018gj and 2020jfo. This observation suggests that weaker Balmer absorption doesn’t necessarily result from a lower mass of the hydrogen envelope in short-plateau SNe. The weaker Balmer absorption is also seen in SNe 2014G and 2020pni in our comparison sample. The spectral features are likely less prominent due to the luminosity from interaction enhancing the ejecta’s temperature (and ionization), leading to an ionization wave penetrating the ejecta inwards from the cold, dense shell (CDS, Chevalier & Fransson, 1994). CDS arises from the higher density of the inner CSM (, see Section 7) resulting from the cooling of shocked areas during the early phases. This phenomenon aligns with observations of Type IIL SNe, which typically exhibit weaker P-Cygni absorption troughs since these SNe likely have more significant interaction with CSM than typical Type IIP SNe (Gutiérrez et al., 2014; Polshaw et al., 2016).
@@ 5.3 Evolution of H : High-Velocity Absorption and Clumpy CSM
The emission peak of has a blueshifted offset by as much as 3,000 at 32 d but evolves towards zero rest velocity by the end of the plateau. This is likely a result of the steep density profile of the ejecta leading to a higher occultation of the receding portion of the ejecta (Anderson et al., 2014b). The blueshifted offset of the emission peak of H (at 30 d) correlates with the decline rate during the plateau phase and the peak luminosity of SNe II (Anderson et al., 2014b). The steep decline rate of SN 2023ixf ( is in agreement with its large offset of 3,000 indicating that the ejecta mass is smaller in SN 2023ixf leading to its shorter plateau. Although this effect is commonly seen in SNe II, this effect is more pronounced in only a handful of events, e.g., SN 2014G (Terreran et al., 2016) and SN 2018gj (Teja et al., 2023b), where the emission peak stays blue-shifted late into the nebular phase.
@@ 5.4 Line velocity Evolution
We compare the velocity evolution of Fe II 5169 for SN 2023ixf with other SNe II in Figure 7. We estimated the line velocity evolution from the blue-shifted absorption trough of their line profiles in the redshift-corrected spectral sequence of SN 2023ixf. We adopt the Fe II 5169 velocities as the photospheric velocity since it forms closest to the photosphere (Dessart & Hillier, 2005) and is the least blended among the iron lines seen in our spectral sequence. The photospheric radius estimated from Fe II absorption trough closely mirrors the blackbody radius as shown in Figure 4. During the mid-plateau phase ( 53 d), the Fe II 5169 velocity inferred for SN 2023ixf is 4350 . This positions it at the higher end of the 1- range when compared with the mean velocity of Fe II 5169 for an extensive collection of SNe II by Gutiérrez et al. (2017a, 3537 851 km/s). This is evident in the spectral comparison during the late-plateau phase (see Figure 5) since we observe higher line blending in SN 2023ixf due to its higher photospheric velocity in comparison to all other SNe II in our comparison except that of SN 2014G.
The trend continues onto the early nebular phase ( 116 d), where SN 2023ixf displays a Fe II 5169 velocity of 3330 compared to the mean value of 2451 679 from Gutiérrez et al. (2017a). The higher photospheric velocity is in agreement with the luminosity-velocity correlation of homologously expanding recombination front of hydrogen (Kasen & Woosley, 2009) since SN 2023ixf is brighter during the mid-plateau phase compared to a normal Type II SN (Anderson et al., 2014a). Additionally, the CSM interaction in SN 2023ixf could also drive the elements at outer/faster regions of the ejecta to be reionized and recombined, leading to a higher estimate of photospheric velocity (Andrews et al., 2019).
@@ 5.5 Early Nebular Phase 90 d
The nebular-phase spectral sequence of SN 2023ixf is shown in the left panel of Figure 8. The nebular spectra of SN 2023ixf displays prominent emission features of H , Mg I] 4571, H , O I 5577, Na I D / He I, [O I] 6300, 6364, H , [Ca II] 7291, 7324 and the Ca II NIR triplet with a flat continuum, typical of SNe II. We observe signs of asymmetry in the line profile of certain emission features during the nebular phase. We observe a dual-peaked axisymmetric888The separation between the two components is less than that of the two components of the O I doublet profile of the [O I] doublet blueshifted by 1500 in SN 2023ixf during the early nebular phase in Figure 9. In addition, we observe an apparent redshifted excess in H and [Ca II] 7291, 7324 at around +5,000 possibly indicating asymmetries in the ejecta. However, the H and [Ca II] showed a single peak symmetric profile with visible signs of redward attenuation as the SN progressed into the nebular phase.
@@
The emergence of an asymmetric emission profile of H and [Ca II] arising due to the attenuation of the red-ward emission from the receding portions of the ejecta is first noticed in our spectral sequence beginning 125 d in Figure 9. This indicates the onset of dust formation and was first noticed in SN 1987A (Lucy et al., 1989). We also observe that the red-blue asymmetry increases as the SN evolves into the nebular phase, indicating an increased dust formation with time (Bevan et al., 2019). The early signatures of dust suggest its formation is happening inside the CDS since the SN ejecta during the 125 – 140 d is too warm for the condensation for molecules (Kozasa et al., 1991). The flash-ionization features in the early spectral sequence and the steep rise in early UV light curves of SN 2023ixf conclusively indicated the presence of a dense CSM. As the shockwave from the SN encountered the denser CSM, it decelerated, compressing the material and increasing the density within the shocked CSM. Radiative cooling then facilitated the emission of photons, aiding in its cooling and forming a distinctive CDS (Chugai, 2009) in the denser CSM. The CDS enables an additional pathway for dust formation in interacting SNe II (Rho et al., 2018). In addition, the clumpiness within the extended CSM encompassing SN 2023ixf facilitates the formation of additional CDS (in addition to its formation in the dense CSM), consequently enhancing molecule formation and eventually forming dust (Inserra et al., 2011). Since nebular phase H arises from the inner ejecta, it wouldn’t show wavelength-dependent attenuation if the dust is formed in the outer CDS. This suggests that the regions where Balmer lines form and dust formation occurs essentially overlap, indicating thorough mixing of the CDS into the inner ejecta following a significant episode of CSM interaction (Bevan et al., 2019).
@@
We also see evidence of flattening in the -band light curve of SN 2023ixf beyond 125 d in Figure 2 evolving at 1.3 0.1 against the relatively consistent decline of 1.8 0.1 in the and bands. This indicates that the continuum luminosity in NIR is evolving steadily; however, the -band light curve is evolving rather slowly due to the emission from CO overtone around 2.3m and/or formation of warm dust. Although we don’t have spectral conformation in NIR, Type II SN 2017eaw (Rho et al., 2018) showed the presence of the first overtone of CO as early as 124 d and also showed flattening in their -band light curves. This strengthens our inference for indirect conformation of molecular CO formation and eventually dust in the case of SN 2023ixf.
@@ 6 Polarimetric Analysis
Figure 10 presents spectropolarimetry observations of SN 2023ixf during the early plateau phase. Due to the low signal-to-noise ratio of the individual epochs, we averaged out the first three epochs (June 14, 15, 16) and the following two epochs (June 18, 19) into two spectropolarimetric observations represented by their average phase of observations. The polarization is expressed as a sum of the and stokes parameters, P , and the polarization angle (PA) is given by the PA = 0.5 .
@@ 6.1 Estimating Interstellar Polarization
To analyze the intrinsic polarization of SN 2023ixf, it’s essential to quantify the Interstellar Polarization (ISP) originating from the irregularly shaped dust particles present across the line of sight in both the Milky Way and the host galaxy M 101. To estimate an upper limit on the ISP, we assume a Serkowski-Galactic type ISP (Serkowski et al., 1975) for M 101 by assuming a similar size and dust composition as the Milky Way. Using the extinction values along the line-of-sight of SN 2023ixf in Section 2, the upper limit on the ISP from Serkowski et al. (1975) is 9 = 0.35%. Since the three peaks in our imaging polarization evolution of SN 2023ixf are all higher than the maximum limit on ISP, we reasonably conclude that the observed polarization characteristics originate from the SN itself.
@@ 6.2 Polarization in the Stokes Q-U Plane and Comparison with other SNe II
In Figure 11, we present ISP-corrected imaging polarization evolution of SN 2023ixf in -band spanning 1.4 d until 140 d. Figure 11 shows 3 distinct peaks in polarization; Point A999We refer to the first observation as point A since we don’t have any data prior to this epoch. [, ] at 1.4 d, Peak B [ at 6.4 d, ] and Peak C [, ] at 79.2 d. The observed stokes parameters ( and ) in -band are shown in Figure 12. Our first imaging polarimetric observation of SN 2023ixf was conducted at 1.4 d, corresponding to Point A, and is associated with the flash-ionized phase due to the interaction of the SN ejecta with an asymmetric, compact, and dense CSM. This is similar to the value of continuum polarization inferred using spectropolarimetric data by Vasylyev et al. (2023). However, we observe a decrease in the polarization signature observed from the dense CSM on the 2nd epoch, i.e., 2.5 d, which is in stark contrast to the rising luminosity (see a six-fold rise in -band luminosity). However, the continuum polarization reported by Vasylyev et al. (2023) at the same epoch, after correction for ISP, indicates a slight polarisation increase from the first epoch, consistent with rising luminosity driven by interaction. The discrepancy could likely be due to the increased line depolarization of H, leading to a decrease in -band polarization as the PA remains the same.
@@
At the next epoch on 4.4 d, we see a change in PA to 55 which later evolves to the polarization peak B at 6.4 d. The polarization peak B lies towards the end of the flash phase transitioning to the early photospheric phase, which shows interaction signatures in the form of intermediate-width P-Cygni profiles (Smith et al., 2023) and close to the emergence of the broad Balmer features at 7 d (see Section 5.2). Polarization at Peak B likely arises from interaction with extended CSM and/or an aspherical shock structure of SN 2023ixf. Following Peak B, we see a decline in polarization across the early plateau phase ( 22 d); however, the PA continues evolving slowly. There is slight disagreement with the continuum PA reported by Vasylyev et al. (2023) likely due to the depolarization of the emission peak of the P-Cygni profile of H. We do observe a revival in the polarization evolution at 28 d with the PA evolving from 100 to 150 . We have a 40-day gap before our next observations at 70 d, before the end of the plateau phase. The observed polarization evolution of SN 2023ixf after 70 d demonstrated a rise towards Peak C in polarization at 79.2 d towards the end of the plateau phase. Since the innermost layers of the ejecta ( 3000 ) are revealed as a result of the entire hydrogen envelope becoming optically thin, this suggests that the electron-scattering atmosphere of the inner core shows a substantial departure from spherical symmetry. The polarization dropped slowly across the early nebular phase, whereas the PA increased gradually. SN 2023ixf shows two preferred dominant axes: a) along the CSM-outer ejecta, where the SN evolved during the early phase until the epoch of polarization minimum on 21.3 d, and b) along the inner ejecta/core, where the SN evolved during the late-plateau and early nebular phase, which is inferred from its polarization evolution.
@@
In a recent study, Nagao et al. (2024) compared the intrinsic polarization properties of 15 SNe II. Broadly, it was inferred that the degree of polarization rises significantly () in the transition/nebular phase; however, it is usually lower during the photospheric phase (). Anomalous cases exist in SN 1987A (Jeffery, 1989) and SN 2021yja (Vasylyev et al., 2024), where a significant polarization is observed during the early plateau phase, gradually declining towards the end of the plateau phase. This is likely driven by electron scattering from interaction with an aspherical CSM, which is analogous to the early continuum polarization (Point A) in SN 2023ixf. However, we also see a secondary peak (Peak B) in polarization in SN 2023ixf, which could possibly arise from interaction with a clumpy low-density extended CSM and/or an aspherical shock structure (see Section 8.4 for further discussions). Towards the end of the plateau phase, the swift increase in polarization during the late-plateau phase and steady decrease during the early nebular phase is similar to that seen in several SNe II like SN 1987A (Jeffery, 1989), SN 2004dj (Leonard et al., 2006) and SN 2013ej (Kumar et al., 2016). Such a polarization evolution is often interpreted as arising from the asymmetries in the inner He-core, and the trend seen is similar in SN 2023ixf. Type IIL SN 2017ahn, which showed early flash-ionization features similar to SN 2023ixf, showed low continuum polarization during the late-plateau phase photospheric and nebular phases, indicating that it might have been asymmetric; however, its viewing angle was lying across its polar axis (Nagao et al., 2021). However, it is clear from the observations of SN 2023ixf that the inclination differs from the axis of symmetry.